# Deep Reinforcement Learning Optimization for Uncertain Nonlinear Systems via Event-Triggered Robust Adaptive Dynamic Programming

### v1

@@ 1 Introduction
Learning-based methods have become a fundamental paradigm in modern engineering systems, enabling algorithms to improve performance through data-driven adaptation without relying solely on explicit mathematical models. Over the past decade, advances in machine learning—particularly in function approximation, optimization, and representation learning—have significantly expanded the capability of intelligent systems operating under uncertainty, compared to traditional analytical methods qin2022real; zhang2024tabnet; hu2025hierarchical. These approaches have been increasingly adopted in control, robotics, and even generative language modelslu2020determination; zhao2024real; tang2025temporal; yao2025learning. However, conventional model-based techniques may be limited in their ability to handle nonlinearities, unknown disturbances, or incomplete system knowledge .
@@
Reinforcement Learning (RL) has gained attention for complex decision-making and control in uncertain, dynamic environments tang2024reinforcement. In control engineering, RL-based methods offer a data-driven alternative to classical model-based designs. This is useful when accurate system models are difficult to obtain. Among these methods, ADP integrates RL with optimal control theory. It facilitates near-optimal control of nonlinear systems by approximating value functions and control policies through function approximators. This eliminates the need to explicitly solve the Hamilton–Jacobi–Bellman (HJB) equation lewis_reinforcement_2009. However, conventional ADP frameworks often rely on continuous or periodic updates to neural network parameters. These updates impose significant computational burdens and may lead to overfitting to transient disturbances or noise.
@@
Event-triggered strategies have been widely adopted in diverse control applications, including networked and embedded systems, multi-agent coordination, and resource-constrained robotic platforms onuoha2024stress; onuoha2024discrete. At the meantime, the ETM has been widely employed in both control and ADP frameworks to reduce computational load (han_event-triggered-based_2024; heemels_introduction_2012; tabuada_event-triggered_2007 ). Unlike time-driven schemes, ETM s update only when systems meet a state- or error-based condition. State deviation or estimation error often directly triggers updates. This approach reduces redundant updates and preserves closed-loop stability (dong2017event; xue_adaptive_2020; onuoha2024stress). By limiting updates to key events, event-triggered ADP boosts efficiency and yields policies less sensitive to disturbances.
@@
Despite these advantages, engineers must ensure robustness against external disturbances and modeling uncertainties. In practice, environmental perturbations, unmodeled dynamics, nonlinear couplings, and parameter uncertainties cause disturbances. Many robust control approaches employ feedback to reduce perturbations rather than explicitly use feedforward compensation (tang2019control; tang2016unmatched; tang2024disturbance). In this context, a ESO estimates the original states and accumulated interference in real time. This allows proactive compensation of parameter mismatches, unmodeled dynamics, and external perturbations in nonlinear systems s(luo2020adaptive; tang2024output; han_PID_ADRC_2009; chen_disturbance-observer-based_2016; ran_new_2021; pu_class_2015; tang2019disturbance).
@@
inspired by the Active Disturbance Rejection Control (ADRC) philosophy (gao2003bandwidth; guo2013MIMO), provide a powerful tool:
@@
Recent work combines ESO-based disturbance rejection with RL for uncertain nonlinear systems (Ran2022Reinforcement; tang2024reinforcementicit). However, these ESORL schemes primarily operate in a time-driven manner: both the controller and learning updates run continuously or periodically, lacking an event-triggered learning mechanism. Many continuous-time ADP designs also impose restrictive Persistence of Excitation (PE) conditions for parameter convergence (jiang2012computational; bian2016actorcritic; kamalapurkar2016SOE), making them hard to verify and enforce in practice.
@@
    A unified control structure incorporating ETM is developed, in which ESO-based state estimation, disturbance compensation, and controller updates occur only at triggering instants. The resulting unified composite control framework enables an aperiodic, computationally efficient implementation of output-feedback RL control. By integrating ESO-based disturbance rejection with event-triggered RL, this work establishes a unified architecture that is not addressed in the existing literature. We propose an event-triggered learning rule for a simulation-of-experience ADP scheme. In this scheme, critic and actor networks update only at triggering instants, using both instantaneous and Extrapolated Bellman Error (EBE). Unlike existing ESORL frameworks (Ran2022Reinforcement; tang2024reinforcementicit), which use continuous or periodic learning, our mechanism yields an aperiodic, data-efficient adaptation. The analysis shows practical stability and Uniform Ultimate Boundedness (UUB) are achieved without a classical PE condition (jiang2012computational; bian2016actorcritic; kamalapurkar2016SOE). It also avoids redundant, high-frequency updates.
@@ 2 Problem Formulation
Following the ADRC philosophy (han_PID_ADRC_2009), the general uncertainty is transferred to a broader state:
@@ 3.1 ESO Design
First, a ESO is designed to predict both the state of the system and all disturbances, following the standard ADRC structure (han_PID_ADRC_2009; chen_disturbance-observer-based_2016):
@@ 3.3 Training Process
To avoid explicitly solving the HJB equation, we adopt an actor–critic architecture enhanced with an ESO-based extrapolation mechanism (Ran2022Reinforcement). The critic network approximates the value function using a linearly parameterized structure
@@ 3.4 Event-Triggered Mechanism
An event is generated only when this inequality is satisfied. Under this rule, all of the closed-loop signals are uniformly bounded, while the number of updates is significantly reduced compared with a periodic scheme (dong2017event).

### v3

@@ 1 Introduction
Learning-based methods have become a fundamental paradigm in modern engineering systems, enabling algorithms to improve performance through data-driven adaptation without relying solely on explicit mathematical models. Over the past decade, advances in machine learning—particularly in function approximation, optimization, and representation learning—have significantly expanded the capability of intelligent systems operating under uncertainty, compared to traditional analytical methods qinRealTimeRemainingUseful2023; zhangTabNetLocallyInterpretable2024; huHierarchicalTestingRabbit2025. These approaches have been increasingly adopted in control, robotics, and even generative language modelsluDeterminationSurfaceCrack2020; zhaoRealTimeObjectDetection2024; tangTemporalScaleTransformer2025; yaoLEARNINGADOPTIONUNDERSTANDING2025. However, conventional model-based techniques may be limited in their ability to handle nonlinearities, unknown disturbances, or incomplete system knowledge .
@@
Reinforcement Learning (RL) has gained attention for complex decision-making and control in uncertain, dynamic environments tangReinforcementLearningBasedApproach2024. In control engineering, RL-based methods offer a data-driven alternative to classical model-based designs. This is useful when accurate system models are difficult to obtain. Among these methods, ADP integrates RL with optimal control theory. It facilitates near-optimal control of nonlinear systems by approximating value functions and control policies through function approximators. This eliminates the need to explicitly solve the Hamilton–Jacobi–Bellman (HJB) equation lewisReinforcementLearningAdaptive2009. However, conventional ADP frameworks often rely on continuous or periodic updates to neural network parameters. These updates impose significant computational burdens and may lead to overfitting to transient disturbances or noise.
@@
Event-triggered strategies have been widely adopted in diverse control applications, including networked and embedded systems, multi-agent coordination, and resource-constrained robotic platforms onuohaStressMatrixBasedFormation2024; onuohaDiscreteTimeStressMatrixBased2024. At the meantime, the ETM has been widely employed in both control and ADP frameworks to reduce computational load (hanEventtriggeredbasedOnlineIntegral2024a; heemelsIntroductionEventtriggeredSelftriggered2012; tabuadaEventTriggeredRealTimeScheduling2007 ). Unlike time-driven schemes, ETM s update only when systems meet a state- or error-based condition. State deviation or estimation error often directly triggers updates. This approach reduces redundant updates and preserves closed-loop stability (dongEventTriggeredAdaptiveDynamic2017b; xueAdaptiveDynamicProgramming2020; onuohaStressMatrixBasedFormation2024). By limiting updates to key events, event-triggered ADP boosts efficiency and yields policies less sensitive to disturbances.
@@
Despite these advantages, engineers must ensure robustness against external disturbances and modeling uncertainties. In practice, environmental perturbations, unmodeled dynamics, nonlinear couplings, and parameter uncertainties cause disturbances. Many robust control approaches employ feedback to reduce perturbations rather than explicitly use feedforward compensation (tang2019control; tangUnmatchedDisturbanceRejection2016; tangDisturbanceObserverBasedOptimal2024). In this context, a ESO estimates the original states and accumulated interference in real time. This allows proactive compensation of parameter mismatches, unmodeled dynamics, and external perturbations in nonlinear systems s(luoAdaptiveAffineFormation2020; tangOutputTrackingUncertain2024; hanPIDActiveDisturbance2009; chenDisturbanceObserverBasedControlRelated2016; ranNewExtendedState2021c; puClassAdaptiveExtended2015; tangDisturbanceRejectionIterative2019).
@@
inspired by the Active Disturbance Rejection Control (ADRC) philosophy (gaoScalingBandwidthparameterizationBased2003; guoConvergenceNonlinearActive2013a), provide a powerful tool:
@@
Recent work combines ESO-based disturbance rejection with RL for uncertain nonlinear systems (ranReinforcementLearningBasedDisturbanceRejection2022; tangReinforcementLearningBasedOutput2024). However, these ESORL schemes primarily operate in a time-driven manner: both the controller and learning updates run continuously or periodically, lacking an event-triggered learning mechanism. Many continuous-time ADP designs also impose restrictive Persistence of Excitation (PE) conditions for parameter convergence (kamalapurkarConcurrentLearningbasedApproximate2016), making them hard to verify and enforce in practice.
@@
    A unified control structure incorporating ETM is developed, in which ESO-based state estimation, disturbance compensation, and controller updates occur only at triggering instants. The resulting unified composite control framework enables an aperiodic, computationally efficient implementation of output-feedback RL control. By integrating ESO-based disturbance rejection with event-triggered RL, this work establishes a unified architecture that is not addressed in the existing literature. We propose an event-triggered learning rule for a simulation-of-experience ADP scheme. In this scheme, critic and actor networks update only at triggering instants, using both instantaneous and Extrapolated Bellman Error (EBE). Unlike existing ESORL frameworks (ranReinforcementLearningBasedDisturbanceRejection2022; tangReinforcementLearningBasedOutput2024), which use continuous or periodic learning, our mechanism yields an aperiodic, data-efficient adaptation. The analysis shows practical stability and Uniform Ultimate Boundedness (UUB) are achieved without a classical PE condition (kamalapurkarConcurrentLearningbasedApproximate2016). It also avoids redundant, high-frequency updates.
@@ 2 Problem Formulation
Following the ADRC philosophy (hanPIDActiveDisturbance2009), the general uncertainty is transferred to a broader state:
@@ 3.1 ESO Design
First, a ESO is designed to predict both the state of the system and all disturbances, following the standard ADRC structure (hanPIDActiveDisturbance2009; chenDisturbanceObserverBasedControlRelated2016):
@@ 3.3 Training Process
To avoid explicitly solving the HJB equation, we adopt an actor–critic architecture enhanced with an ESO-based extrapolation mechanism (ranReinforcementLearningBasedDisturbanceRejection2022). The critic network approximates the value function using a linearly parameterized structure
@@ 3.4 Event-Triggered Mechanism
An event is generated only when this inequality is satisfied. Under this rule, all of the closed-loop signals are uniformly bounded, while the number of updates is significantly reduced compared with a periodic scheme (dongEventTriggeredAdaptiveDynamic2017b).